April needs to wrap a box that is 3 feet long, 2 feet wide and 1 foot high. How much paper will she need?
Answer by Boreal(15235)

The box has 6 parts:
2 lengths and widths (top and bottom)
2 lengths and heights (sides)
2 widths and heights (ends)
The first is 2 * 6 ft^2=12 ft^2
The second is 2*3 ft^2=6 ft^2 (the height is 1)
The third is 2*2 ft^2=4 ft^2 (height is 1)
22 ft^2